OpenStudy (anonymous):

Solve 7x = –21. A. –14 B. –3 C. –28 D. –21

OpenStudy (anonymous):

divide both sides by 7

OpenStudy (anonymous):

x=-3 is your answer

OpenStudy (shamil98):

7x = –21. 7x/7 = -21/7 x = -3